### v2

@@ 5.2 The infinite-dimensional learning problem and its fidelity
Let the coefficients $(\widehat{a},\widehat{b})\in\mathcal{H}^{+}_{m,R_{*}}$ be the unique solution to LP with ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$. With probability at least $1-\delta$, it holds that:
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely.
@@
$\underset{t\in[0,T]}{\sup}\ \|p_{\widehat{a},\widehat{b}}(t,\cdot)-\widehat{p}(t,\cdot)\|^{2}_{L^{2}}\leq C\big{(}\|(\widehat{a},\widehat{b})\|_{\mathcal{H}_{m}}\big{)}C(a_{*},b_{*}){\color[rgb]{0,0,0}\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}},$
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely, whereas the constant $C\big{(}\|(\widehat{a},\widehat{b})\|_{\mathcal{H}_{m}}\big{)}>0$ continuously depends on $\|(\widehat{a},\widehat{b})\|_{\mathcal{H}_{m}}$ uniquely. Up to overloading these constants, combined with (17) this latter inequality readily yields
(19) $\int^{T}_{0}\|p_{\widehat{a},\widehat{b}}(t,\cdot)-p(t,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}t\leq C\big{(}\|(\widehat{a},\widehat{b})\|_{\mathcal{H}_{m}}\big{)}C(a_{*},b_{*}){\color[rgb]{0,0,0}\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}.$
@@
(20) $\|(\widehat{a},\widehat{b})\|^{2}_{\mathcal{H}_{m}}\leq\frac{C(a_{*},b_{*})}{\lambda}\left(\lambda+{\color[rgb]{0,0,0}\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}\right)=2C(a_{*},b_{*}),$
@@ 5.3 The finite-dimensional learning problem and its fidelity
We start by recalling and adapting the approximation tools we introduced in Section 4 to our framework. Let $D=[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}$, and consider the RKHS $\mathcal{H}_{D}=H^{d(m)}(D,\mathbb{R})$ with associated kernel $K_{D}$. Without loss of generality, we may assume $K_{D}$ and all its derivatives equal zero on $[0,T]\times\partial B^{\mathbb{R}^{n}}_{R_{*}}(0)$. For any set of $Q\in\mathbb{N}$ points
@@
$\begin{split}\tilde{a}(t,y)&\triangleq\sum_{\ell,\ell^{\prime}=1}^{Q}A_{\ell,\ell^{\prime}}\,K_{D}((t,y),(t_{\ell},x_{\ell}))\,K_{D}((t,y),(t_{\ell^{\prime}},x_{\ell^{\prime}})),\\ &\hskip 172.22168pt(t,y)\in D,\quad i,j\in\{1,\dots,n\},\\ \tilde{b}(t,x)&\triangleq\sum_{\ell=1}^{Q}B_{\ell}K_{D}((t,y),(t_{\ell},x_{\ell})),\quad(t,y)\in D,\quad i\in\{1,\dots,n\},\end{split}$
@@
(21) $\tilde{a}(t,y)=\Phi_{Q}(t,y)A\Phi_{Q}(t,y)^{\top},\quad\tilde{b}(t,y)=\Phi_{Q}(t,y)B,$
with $B\triangleq(B_{1},\dots,B_{Q})^{\top}\in\mathbb{R}^{Qn}$, and $A\in\mathbb{R}^{Qn\times Qn}$ is the $Q\times Q$ block matrix with block entries $A_{\ell,\ell^{\prime}}$. Therefore, we define the finite dimensional convex subset of $\mathcal{H}^{+}_{m,R_{*}}$ in which the tuple $(\widehat{a},\widehat{b})$ is approximated to be (zero is imposed as value outside $D$)
$\mathcal{H}^{+,Q}_{m,R_{*}}\triangleq\textnormal{span}\ \Big{\{}(\tilde{a},\tilde{b}):\ A\in\mathbb{R}^{Qn\times Qn},\ A\succeq 0\ \textnormal{and}\ B\in\mathbb{R}^{Qn}\Big{\}}\subseteq\mathcal{H}^{+}_{m,R_{*}}.$
Before moving to the core of this section, by leveraging the facts we recalled in Section 4, we provide a crucial approximation results for coefficients $(a,b)\in\mathcal{H}^{+}_{m,R_{*}}$. Specifically, by combining the bounds for Sobolev functions with scattered zeros we listed in Section 4 with Theorem 2.1, we obtain the following:
@@
There exists a constant $C>0$ such that, for every $(a,b)\in\mathcal{H}^{+}_{m,R_{*}}$ with $a(t,y)\succ\beta I$, $(t,y)\in D$, for some constant $\beta>0$, there exists a tuple $(\mathcal{P}_{Q}(a),\mathcal{P}_{Q}(b))\in\mathcal{H}^{+,Q}_{m,R_{*}}$:
$\|(a,b)-(\mathcal{P}_{Q}(a),\mathcal{P}_{Q}(b))\|_{W^{2,\infty}}\leq C(\|\sqrt{a}\|^{2}_{H^{d(m)}}+\|b\|_{H^{d(m)}})h^{2m-1}_{Q},$
$\|(\mathcal{P}_{Q}(a),\mathcal{P}_{Q}(b))\|_{\mathcal{H}_{m}}\leq C(\|\sqrt{a}\|^{2}_{H^{d(m)}}+\|b\|_{H^{d(m)}}).$
@@
First, by combining the bounds for Sobolev functions with scattered zeros Eq. 13 in Section 4 with Theorem 2.1, we may readily claim the existence of a constant $C>0$ (implicitly overloaded below) and $\mathcal{P}_{Q}(b)\triangleq\tilde{b}$ in the form Eq. 21 such that
$\|b-\tilde{b}\|_{W^{2,\infty}}\leq C_{1}\|b\|_{H^{d(m)}}h^{2m-1}_{Q}.$
We now turn to the construction of $\mathcal{P}_{Q}(a)\triangleq\tilde{a}$. This is slightly more elaborated and is taken from [40, 31, 39, 29], in particular, adapting the construction of [31]. Note that the matrix square root $\sqrt{\cdot}$ has uniformly bounded derivatives of any order on the set $\{A\in\mathbb{R}^{n\times n}|A\succ\beta I\}$. By assumption, it thus follows that $\sqrt{a}\in H^{d(m)}(D,\mathbb{R}^{n\times n})$, and in particular that the element-wise functions $(\sqrt{a})_{i,j}\triangleq e_{i}\ \sqrt{a}\ e_{j}\in H^{d(m)}(D,\mathbb{R})$. Now, we can build our approximation $\tilde{a}$ as follows
$\tilde{a}(t,x)\triangleq v(t,y)^{\top}v(t,y)\succeq 0,\quad(t,y)\in D,$
where $v:D\to\mathbb{R}^{n\times n}$ is obtained by element-wise applying the bound for Sobolev functions with scattered zeros Eq. 13 in Section 4. In particular, we obtain that
$\|(\sqrt{a})_{i,j}-v_{i,j}\|_{W^{2,\infty}}\leq C\|\sqrt{a}_{i,j}\|_{H^{d(m)}}h^{2m-1}_{Q},\quad i,j\in\{1,\dots,n\}.$
Therefore, since $\sqrt{a}^{\top}\sqrt{a}-v^{\top}v=\sqrt{a}^{\top}(\sqrt{a}-v)+(\sqrt{a}-v)^{\top}\sqrt{a}-(\sqrt{a}-v)^{\top}(\sqrt{a}-v)$, and $W^{2,\infty}$ is in particular a Banach algebra, we may infer that
$\|a-\tilde{a}\|_{W^{2,\infty}}\leq 2\|\sqrt{a}-v\|_{W^{2,\infty}}\|\sqrt{a}\|_{W^{2,\infty}}+\|\sqrt{a}-v\|_{W^{2,\infty}}^{2}.$
At this step, the inequalities
$\|\sqrt{a}-v\|^{2}_{W^{2,\infty}}\leq\sum_{i,j=1}^{n}\|(\sqrt{a})_{ij}-v_{ij}\|^{2}_{W^{2,\infty}}$
and $\|\cdot\|_{W^{2,\infty}}\leq C\|\cdot\|_{H^{d(m)}}$ may be combined to readily derive that
$\|a-\tilde{a}\|_{W^{2,\infty}}\leq Cn^{2}\|\sqrt{a}\|_{H^{d(m)}}^{2}h^{2m-1}_{Q},$
that is the first inequality claimed in the theorem. In particular, note that
@@
for some $R_{1},\dots,R_{Q}\in\mathbb{R}^{n\times n}$. By denoting $R=(R_{1},\dots,R_{Q})\in\mathbb{R}^{n\times Qn}$, we have that
$\tilde{a}(t,y)=\Phi_{Q}(t,y)A\Phi_{Q}(t,y)^{\top},\quad(t,y)\in D,$
with $A\triangleq RR^{\top}\in\mathbb{R}^{Qn\times Qn}$ semi-positive-definite. We thus infer that $(\tilde{a},\tilde{b})\in\mathcal{H}^{+,Q}_{m,R_{*}}$.
To conclude, according to the results we gathered in Section 4, it holds that $\tilde{b}=\Pi_{Q}(b)$ and $v=\Pi_{Q}(\sqrt{a})$, where $\Pi_{Q}$ denotes the corresponding projection operator of each involved RKHS, and thus $\tilde{a}=\Pi_{Q}(\sqrt{a})^{\top}\Pi_{Q}(\sqrt{a})$. This implies that
$\|\Pi_{Q}(b)\|_{H^{d(m)}}\leq\|b\|_{H^{d(m)}}\quad\textnormal{and}\quad\|\Pi_{Q}(\sqrt{a})\|_{H^{d(m)}}\leq\|\sqrt{a}\|_{H^{d(m)}}.$
The proof is concluded considering that $H^{d(m)}$ is in particular a Banach algebra.
In the rest of the manuscript, we adopt the notation we introduced in Theorem 5.12. We are now ready to define our finite-dimensional problem for learning the coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m,R_{*}}$, which writes as follows:
@@
Since $\mathcal{H}^{+,Q}_{m,R_{*}}$ is a finite-dimensional convex set, the well-posedness of LPQ may be proven similarly to the well-posedness of the original problem LP, i.e., by replicating the proof of Proposition 5.6. We thus report this result below without proof:
@@
For every $\lambda>0$, problem LPQ is well-posed and has a unique solution, which is denoted by $(\widehat{a}_{Q},\widehat{b}_{Q})\in\mathcal{H}^{+,Q}_{m,R_{*}}$.
Remark 5.16.
Thanks to Assumption $(A)$, up to re-scaling $\alpha$ and overloading $a_{*}$, one may always assume that $a_{*}(t,y)\succ\alpha I$, $(t,y)\in D$. This property would enable applying Theorem 5.12 directly to $(a_{*},b_{*})$. Unfortunately, the setting $a_{*}(t,y)\succ\alpha I$, $(t,y)\in D$, does not fit Assumption $(B)$, and in turn the problem formulation of LPQ. Nevertheless, one may appropriately tailor the definition of $\mathcal{H}^{+,Q}_{m,R_{*}}$ to allow for $a_{*}(t,y)\succ\alpha I$, $(t,y)\in D$. Indeed, in such setting Assumption $(B)$ would rather require that $\textnormal{supp}\big{(}a(t,\cdot)-\alpha I,b(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}$, for every $t\in[0,T]$. Therefore, we should consider a larger domain $D$ in which learning is pursued, e.g., $D\triangleq[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}}$, selecting $K_{D}$ to equal 1 on $[0,T]\times\partial B^{\mathbb{R}^{n}}_{R_{*}+1}(0)$ and such that all its derivatives equal zero on $[0,T]\times\partial B^{\mathbb{R}^{n}}_{R_{*}+1}(0)$ (this selection is always possible under mild assumptions). Thanks to this choice, $\mathcal{H}^{+,Q}_{m,R_{*}}$ may be replaced by the following subset
$\textnormal{span}\ \Big{\{}(\chi\tilde{a}+(1-\chi)\alpha I,\tilde{b}):$ $A\in\mathbb{R}^{Qn\times Qn},\ A\succeq\alpha I,\ \textnormal{and}\ B\in\mathbb{R}^{Qn}\Big{\}}\subseteq\mathcal{H}^{+}_{m,R_{*}+1},$
where the definition of $\mathcal{H}^{+}_{m,R_{*}+1}$ corresponds to $\mathcal{H}^{+}_{m,R_{*}}$’s with updated requirement $\textnormal{supp}\big{(}a(t,\cdot)-\alpha I,b(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}$. Here, $\chi\in C^{\infty}([0,T]\times\mathbb{R}^{n},[0,1])$ is some fixed cut-off function satisfying $\chi|_{\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}}=1$ and $\textnormal{supp}\;\chi\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}$. One readily check this updated definition does not affect the validity of Theorem 5.12, and in turn of the rates of convergence we will develop shortly. From a numerical point of view, the computational cost of solving LPQ with this updated definition of $\mathcal{H}^{+,Q}_{m,R_{*}}$ remains unchanged (see Section 6). Therefore, without loss of generality, from now on we assume we may apply Theorem 5.12 to $(a_{*},b_{*})\in\mathcal{H}^{+}_{m,R_{*}}$ without further ado.
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely.
Proof 5.18.
@@
$L_{\lambda}(\widehat{a}_{Q},\widehat{b}_{Q})-\lambda\|\big{(}\mathcal{P}_{Q}(a_{*}),\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq$
@@
By combining this inequality with Theorem 5.12 (and Remark 5.16), we infer that
@@
(22) $L_{\lambda}(\widehat{a}_{Q},\widehat{b}_{Q})-\lambda\|\big{(}\mathcal{P}_{Q}(a_{*}),\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq$

### v3

@@ 5.2 The infinite-dimensional learning problem and its fidelity
Let the coefficients $(\widehat{a},\widehat{b})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$ be the unique solution to LP with ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$. With probability at least $1-\delta$, it holds that:
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely (in particular, $C(a_{*},b_{*})$ implicitly depends on the hyper-parameters $m$, $p_{0}$, and $\alpha$).
@@
$\underset{t\in[0,T]}{\sup}\ \|p_{\widehat{a},\widehat{b}}(t,\cdot)-\widehat{p}(t,\cdot)\|^{2}_{L^{2}}\leq C\big{(}\alpha,\|(\widehat{a}-\alpha I,\widehat{b})\|_{\mathcal{H}_{m}}\big{)}C(a_{*},b_{*})\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2},$
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely, whereas the constant ${\color[rgb]{0,0,0}C\big{(}\alpha,\|(\widehat{a}-\alpha I,\widehat{b})\|_{\mathcal{H}_{m}}\big{)}}$ continuously depends on ${\color[rgb]{0,0,0}\alpha}$ and ${\color[rgb]{0,0,0}\|(\widehat{a}-\alpha I,\widehat{b})\|_{\mathcal{H}_{m}}}$ uniquely. Up to overloading these constants, combined with (17) this latter inequality readily yields
(19) $\int^{T}_{0}\|p_{\widehat{a},\widehat{b}}(t,\cdot)-p(t,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}t\leq C\big{(}\alpha,\|(\widehat{a}-\alpha I,\widehat{b})\|_{\mathcal{H}_{m}}\big{)}C(a_{*},b_{*})\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}.$
@@
(20) $\|(\widehat{a}-\alpha I,\widehat{b})\|^{2}_{\mathcal{H}_{m}}\leq\frac{C(a_{*},b_{*})}{\lambda}\left(\lambda+\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}\right)=2C(a_{*},b_{*}),$
@@ 5.3 The finite-dimensional learning problem and its fidelity
We start by recalling and adapting the approximation tools we introduced in Section 4 to our framework. Let ${\color[rgb]{0,0,0}D=[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}}$, and consider the RKHS $\mathcal{H}_{D}=H^{d(m)}(D,\mathbb{R})$ with associated kernel $K_{D}$. Without loss of generality, we may assume $K_{D}$ to equal 1 on $[0,T]\times\partial B^{\mathbb{R}^{n}}_{R_{*}+1}(0)$ and such that all its derivatives equal zero on $[0,T]\times\partial B^{\mathbb{R}^{n}}_{R_{*}+1}(0)$. While not limiting, pursuing learning with slightly larger domains, i.e., $B^{\mathbb{R}^{n}}_{R_{*}+1}(0)$ instead of $B^{\mathbb{R}^{n}}_{R_{*}}(0)$, is key to enable rates of convergence under positiveness constraints, as we will see shortly. For any set of $Q\in\mathbb{N}$ points
@@
$\begin{split}{\color[rgb]{0,0,0}a_{A}}(t,y)&\triangleq\sum_{\ell,\ell^{\prime}=1}^{Q}A_{\ell,\ell^{\prime}}\,K_{D}((t,y),(t_{\ell},x_{\ell}))\,K_{D}((t,y),(t_{\ell^{\prime}},x_{\ell^{\prime}})),\\ &\hskip 172.22168pt(t,y)\in D,\quad i,j\in\{1,\dots,n\},\\ {\color[rgb]{0,0,0}b_{B}}(t,x)&\triangleq\sum_{\ell=1}^{Q}B_{\ell}K_{D}((t,y),(t_{\ell},x_{\ell})),\quad(t,y)\in D,\quad i\in\{1,\dots,n\},\end{split}$
@@
(21) ${\color[rgb]{0,0,0}a_{A}}(t,y)=\Phi_{Q}(t,y)A\Phi_{Q}(t,y)^{\top},\quad{\color[rgb]{0,0,0}b_{B}}(t,y)=\Phi_{Q}(t,y)B,$
with $B\triangleq(B_{1},\dots,B_{Q})^{\top}\in\mathbb{R}^{Qn}$, and $A\in\mathbb{R}^{Qn\times Qn}$ is the $Q\times Q$ block matrix with block entries $A_{\ell,\ell^{\prime}}$. Therefore, we define the finite dimensional convex subset in which $(\widehat{a},\widehat{b})$ is approximated by
${\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}\triangleq\Bigg{\{}\bigg{(}\chi\left(a_{A}+\frac{\alpha}{2}I\right)+(1-\chi)\alpha I,\chi b_{B}\bigg{)}:}$ ${\color[rgb]{0,0,0}A\in\mathbb{R}^{Qn\times Qn},}$
${\color[rgb]{0,0,0}A\succeq 0,\ \textnormal{and}\ B\in\mathbb{R}^{Qn}\Bigg{\}}\subseteq{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha/2}_{m}},}$
where $\chi\in C^{\infty}(\mathbb{R}^{n},[0,1])$ is any given cut-off function satisfying $\chi|_{\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}}=1$ and $\textnormal{supp}\;\chi\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}$. Through the subset ${\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$, we enlarge the space in which we look for finite-dimensional approximating coefficients, due to $a_{A}+\alpha/2$ with $A\succeq 0$. Crucial to enable rates of convergence under positiveness constraint as we will see shortly, such strategy is not limiting, as it poorly affects computational complexity.
Before moving to the core of this section, by leveraging the facts we recalled in Section 4, we provide a crucial approximation results for coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$. Specifically, by combining the bounds for Sobolev functions with scattered zeros we listed in Section 4 with Theorem 2.1, we obtain the following:
@@
There exists a constant $C>0$ such that, for every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$, there exists a tuple $(\mathcal{P}_{Q}(a),\mathcal{P}_{Q}(b))\in{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$ satisfying:
$\|(a,b)-(\mathcal{P}_{Q}(a),\mathcal{P}_{Q}(b))\|_{W^{2,\infty}}\leq C(\|\sqrt{a}\|^{2}_{H^{d(m)}(D)}+\|b\|_{H^{d(m)}})h^{2m-1}_{Q},$
$\|(\mathcal{P}_{Q}(a)-\alpha I,\mathcal{P}_{Q}(b))\|_{\mathcal{H}_{m}}\leq C(\|\sqrt{a}\|^{2}_{H^{d(m)}(D)}+\|b\|_{H^{d(m)}}).$
@@
Recall that, by definition, $a(t,\cdot)=\alpha\,I$ and $b(t,\cdot)=0$ outside of $\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}$. First, thanks to the bounds for Sobolev functions with scattered zeros Eq. 13, we may readily claim the existence of a constant $C>0$ (implicitly overloaded below) and of a map $b_{B}$ in the form Eq. 21 such that
$\|b-b_{B}\|_{H^{d(m)-2m+1}(D)}\leq C\|b\|_{H^{d(m)}}h^{2m-1}_{Q}.$
Therefore, on the one hand we may compute
$\|b_{B}\|_{H^{d(m)-2m+1}\big{(}[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}\setminus\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}\big{)}}=$
$=\|b-b_{B}\|_{H^{d(m)-2m+1}\big{(}[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}\setminus\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}\big{)}}\leq C\|b\|_{H^{d(m)}}h^{2m-1}_{Q},$
and on the other, combining the two previous inequalities yields
$\|b-\chi b_{B}\|_{H^{d(m)-2m+1}}=\|b-\chi b_{B}\|_{H^{d(m)-2m+1}(D)}$
$\leq\|b-b_{B}\|_{H^{d(m)-2m+1}\big{(}[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}\big{)}}+\|\chi b_{B}\|_{H^{d(m)-2m+1}\big{(}[0,T]\times\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}\setminus\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}\big{)}}$
$\leq C\|b\|_{H^{d(m)}}h^{2m-1}_{Q}.$
In particular, leveraging Theorem 2.1 we may thus select $\mathcal{P}_{Q}(b)\triangleq\chi b_{B}$.
We now turn to the construction of the approximation $\mathcal{P}_{Q}(a)$. This construction is slightly more elaborated and is taken from [40, 31, 39, 29], in particular, adapting the arguments in [31]. Note that the matrix square root $\sqrt{\cdot}$ has uniformly bounded derivatives of any order on the set $\{A\in\mathbb{R}^{n\times n}:\ A\succeq\alpha/2\,I\}$. By assumption, it in particular follows that $\sqrt{a-\alpha/2\,I}\in H^{d(m)}(D,\mathbb{R}^{n\times n})$ (we are implicitly restricting the map $a$ to the domain $D$), and therefore that $(\sqrt{a})_{ij}\triangleq e_{i}\ \sqrt{a-\alpha/2\,I}\ e_{j}\in H^{d(m)}(D,\mathbb{R})$. Now, we can build our approximation $a_{A}$ as follows:
$a_{A}(t,x)\triangleq v(t,y)^{\top}v(t,y),\quad(t,y)\in D,$
where $v:D\to\mathbb{R}^{n\times n}$ is obtained by element-wise applying the bound for Sobolev functions with scattered zeros Eq. 13 to $\sqrt{a-\alpha/2\,I}$, which yields
(22) $\left\|\left(\sqrt{a-\frac{\alpha}{2}I}\right)_{ij}-v_{ij}\right\|_{H^{d(m)-2m+1}(D)}\leq$
$\leq C\left\|\left(\sqrt{a-\frac{\alpha}{2}I}\right)_{ij}\right\|_{H^{d(m)}(D)}h^{2m-1}_{Q}\leq C\|\sqrt{a}\|_{H^{d(m)}(D)}h^{2m-1}_{Q},\quad i,j\in\{1,\dots,n\}.$
In particular, since
@@
for some $R_{1},\dots,R_{Q}\in\mathbb{R}^{n\times n}$, by denoting $R=(R_{1},\dots,R_{Q})\in\mathbb{R}^{n\times Qn}$, we have that
(23) $a_{A}(t,y)=\Phi_{Q}(t,y)A\Phi_{Q}(t,y)^{\top},\quad(t,y)\in D,$
with $A\triangleq RR^{\top}\in\mathbb{R}^{Qn\times Qn}$ and $A\succeq 0$. At this step, thanks to bound (22), the identity $\sqrt{M}^{\top}\sqrt{M}-v^{\top}v=\sqrt{M}^{\top}(\sqrt{M}-v)+(\sqrt{M}-v)^{\top}\sqrt{M}-(\sqrt{M}-v)^{\top}(\sqrt{M}-v)$, for $M\in\mathbb{R}^{n\times n}$, and the fact that $H^{d(m)-2m+1}(D)$ is a Banach algebra, we first infer that
$\bigg{\|}\left(a-\frac{\alpha}{2}I\right)-$ $a_{A}\bigg{\|}_{H^{d(m)-2m+1}(D)}\leq\left\|\sqrt{a-\frac{\alpha}{2}I}-v\right\|_{H^{d(m)-2m+1}(D)}^{2}$
$+2\left\|\sqrt{a-\frac{\alpha}{2}I}-v\right\|_{H^{d(m)-2m+1}(D)}\left\|\sqrt{a-\frac{\alpha}{2}I}\right\|_{H^{d(m)-2m+1}(D)}.$
Then, combining this latter inequality with
$\left\|\sqrt{a-\frac{\alpha}{2}I}-v\right\|^{2}_{H^{d(m)-2m+1}(D)}\leq\sum_{i,j=1}^{n}\left\|\left(\sqrt{a-\frac{\alpha}{2}I}\right)_{ij}-v_{ij}\right\|^{2}_{H^{d(m)-2m+1}(D)},$
identity (22), and Theorem 2.1 (readily extended to bounded domains) yields
$\left\|\left(a-\frac{\alpha}{2}I\right)-a_{A}\right\|_{W^{2,\infty}(D)}\leq$
$\leq C\left\|\left(a-\frac{\alpha}{2}I\right)-a_{A}\right\|_{H^{d(m)-2m+1}(D)}\leq C\|\sqrt{a}\|_{H^{d(m)}(D)}^{2}h^{2m-1}_{Q}.$
Thanks to this latter inequality, we may therefore compute
$\left\|\chi\left(a_{A}+\frac{\alpha}{2}I\right)+(1-\chi)\alpha I-a\right\|_{W^{2,\infty}}\leq$
$\leq\left\|\chi\left(a_{A}+\frac{\alpha}{2}I\right)+(1-\chi)\alpha I-a\right\|_{W^{2,\infty}(D)}$
$\hskip 129.16626pt+\|\alpha I-a\|_{W^{2,\infty}\big{(}[0,T]\times\mathbb{R}^{n}\setminus\overline{B^{\mathbb{R}^{n}}_{R_{*}+1}(0)}\big{)}}$
$\leq\left\|\chi\left(\left(a_{A}+\frac{\alpha}{2}I\right)-a\right)\right\|_{W^{2,\infty}(D)}+\|(1-\chi)(\alpha I-a)\|_{W^{2,\infty}(D)}$
$\leq C\left\|\left(a-\frac{\alpha}{2}I\right)-a_{A}\right\|_{W^{2,\infty}(D)}+C\|\alpha I-a\|_{W^{2,\infty}\big{(}[0,T]\times\mathbb{R}^{n}\setminus\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)}\big{)}}$
$\leq C\|\sqrt{a}\|_{H^{d(m)}(D)}^{2}h^{2m-1}_{Q},$
so that we may in conclusion select $\mathcal{P}_{Q}(a)\triangleq\chi(a_{A}+\alpha/2\,I)+(1-\chi)\alpha I$.
To conclude note that, according to the results in Section 4, it holds that $b_{B}=\Pi_{Q}(b)$ and $v=\Pi_{Q}(\sqrt{a-\alpha/2\,I})$, where $\Pi_{Q}$ denotes the corresponding projection operator of each involved RKHS, and thus $a_{A}=\Pi_{Q}(\sqrt{a-\alpha/2\,I})^{\top}\Pi_{Q}(\sqrt{a-\alpha/2\,I})$. This implies that
$\|\Pi_{Q}(b)\|_{H^{d(m)}(D)}\leq$ $\|b\|_{H^{d(m)}(D)}\quad\textnormal{and}$
$\|\Pi_{Q}(\sqrt{a-\alpha/2\,I})\|_{H^{d(m)}(D)}\leq\|\sqrt{a-\alpha/2\,I}\|_{H^{d(m)}(D)}.$
The proof readily follows from these latter inequalities, considering that
$\|\mathcal{P}_{Q}(a)-\alpha I\|_{H^{d(m)}}\leq C\left\|a_{A}-\frac{\alpha}{2}I\right\|_{H^{d(m)}(D)},$
and that $H^{d(m)}(D)$ is in particular a Banach algebra.
In the rest of the manuscript, we adopt the notation we introduced in Theorem 5.12. We are now ready to define our finite-dimensional problem for learning the coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$, which writes as follows:
@@
Since ${\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}{\color[rgb]{0,0,0}-\,\{(\alpha I,0)\}\subseteq\mathcal{H}_{m}}$ is closed and convex, the well-posedness of LPQ may be proven similarly to the well-posedness of the problem LP, i.e., by replicating the proof of Proposition 5.6. We thus report this result below without proof:
@@
For every $\lambda>0$, problem LPQ is well-posed and has a unique solution, which is denoted by $(\widehat{a}_{Q},\widehat{b}_{Q})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$.
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely (in particular, $C(a_{*},b_{*})$ implicitly depends on the hyper-parameters $m$, $p_{0}$, and $\alpha$).
Proof 5.17.
@@
$L_{\lambda}(\widehat{a}_{Q},\widehat{b}_{Q})-\lambda\|\big{(}\mathcal{P}_{Q}(a_{*})-\alpha I,\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq$
@@
By combining this inequality with Theorem 5.12, we infer that
@@
(24) $L_{\lambda}(\widehat{a}_{Q},\widehat{b}_{Q})-\lambda\|\big{(}\mathcal{P}_{Q}(a_{*})-\alpha I,\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq$